OpenStudy (anonymous):

Convert the parametric equation into polar form:

OpenStudy (anonymous):

\[x = \frac{ 3t }{ 1+t^{3} }\] \[y = \frac{ 3t^{2} }{ 1+t^{3} }\]

OpenStudy (luigi0210):

Do you know the formula for converting from parametric to polar and the other way around?

OpenStudy (anonymous):

Wasn't even provided a formula. I'm used to eliminating the parameter normally or doing simple rectangular equations to polar, but this is kinda smack in your face brand new, so not even sure where to start.

OpenStudy (luigi0210):

@Loser66 Feel free to jump in bud.

OpenStudy (luigi0210):

Probably should of asked this earlier, but what class is this for?

OpenStudy (anonymous):

Well, this is my own kind of personal review and skimming through stuff, but this is end calc 2 intro calc 3.

OpenStudy (luigi0210):

@Jonask can probably help you better then, I am not into calculus 3 or even 2 yet.

OpenStudy (anonymous):

You were thinking regular rectangular to polar and vice versa @Luigi0210 ? :P

OpenStudy (luigi0210):

Yea, exactly xD

OpenStudy (anonymous):

\[x=\frac{3t}{1+t^3},y=\frac{3t^2}{1+t^3}\] \[y=t\frac{3t}{1+t^3}=tx\] so \[t=\frac{y}{x}\] then \[x=\frac{3t}{1+t^3}=\frac{\frac{y}{x}}{1+(\frac{y}{x})^3}=x\] \[\frac{yx^3}{(x^3+y^3)x}=x\\x^2(x^3+y^3)=yx^3\\x^3+y^3=yx\]

OpenStudy (anonymous):

Forgot a 3 in there, but yeah, I see it. I guess it's just being more clever and better at recognition than what I am, lol.

OpenStudy (anonymous):

i think we can use polar coordintes wen we eliminated t @Loser66

OpenStudy (anonymous):

the 3 doesnt have to be there notice that its part of x @Concentrationalizing

OpenStudy (anonymous):

Is it? O.o It just looks like it disappeared once you replaced t with y/x

OpenStudy (anonymous):

@Loser66 has a good substitution,also you ight jus substitute in the last equation to get \[\cos^3\theta+\sin^3\theta=\cos\theta\sin\theta\]

OpenStudy (anonymous):

no its not about clever,we are jus trying to eliminate t by making the two equations agree,the long way will be solving for t then substituting but recognising the combination is better in this case

OpenStudy (anonymous):

@Loser66 it makes lots of sense

OpenStudy (anonymous):

Exactly, recognizing it as a combination. I justimmediately see that it'd be hardtotake one of the parameters and solve for t. And then some recent problems had it where you would go from polar to parametric, so I was trying to think of that but backwards, except that didn't work. So....I suppose recent problems and what I'm used to just limited my thinking. So yes, I still say recognition, lol.

OpenStudy (anonymous):

Yeah, and I never was able to figure out my last problem with the series solution @Loser66 haha

OpenStudy (anonymous):

Well, from where you had it before was just fine. Although I still see a 3 needed in there, lol. \[x^{3} + y^{3} = 3xy\] \[r^{3}\cos^{3}\theta + r^{3}\sin^{3}\theta = 3rcos \theta r \sin \theta\] \[r*r^{2}(\cos^{3}\theta + \sin^{3} \theta) = 3r^{2}\cos \theta \sin \theta\] \[r = \frac{ 3\cos \theta \sin \theta }{ \cos^{3} \theta + \sin^{3} \theta}\] dun dun dun

OpenStudy (anonymous):

yes you are on the spot but still i dont c the 3 why it has to be here haha

OpenStudy (anonymous):

\[r=\frac{\sin \theta \cos\theta}{\sin^3 \theta +\cos^3\theta}\]

OpenStudy (anonymous):

\[x = \frac{ 3t }{ 1+t^{3}} \] \[t = y/x\] \[x = \frac{ \frac{ 3y }{ x } }{ 1+(\frac{ y }{ x })^{3} }\] 3 still would be there?

OpenStudy (anonymous):

i see it...thank you,you r right

OpenStudy (anonymous):

Alrighty ^_^ Lol, and yeah, luigi didnt realize it was cal2 either @Loser66 lol. But yeah, thanks everyone : )

OpenStudy (anonymous):

Im all over the place with the problems I post. But most of the time expect cal 2, 3, diff eq, linear algebra or something xD

OpenStudy (anonymous):

Im pretty much done with all my calc2 review after one section, then Im jumping back to finish off the series solutions to DEs and then calc 3 we go.

OpenStudy (anonymous):

The first DE course I just did was easy. But when I did calc 2, we never quite touched on this last section with parametrics and polar and such. A lot of it is trig review, so np, but some of this is new. So I'm basically finishing off what I shouldve done before and the last chapter in DE we didn't finish. Once I'm done with that it's full-board calc 3 and linear algebra.

OpenStudy (anonymous):

lool i jus started DE n calc 2...u can crawl easier wen u watch ocw lectures,the material is great

OpenStudy (anonymous):

I'm a math major, but I also do language study as a hobby and personal interest. Oh, special circumstances. Normally you can't take it before multivariable, but the reason I can is it's more of a education war between the community college here and the main university, at least it seems that way. The community college offers some of the higher level courses as a way to kinda steal from the university I suppose. If I took the DE course at uni it'd be above linear, multivariable, all of that.

OpenStudy (anonymous):

Of course taking it at the community college BEFORE multivariable and linear has its minor challenges, but the calc 3 and linear portions of the course I hadto do were np for me to catch up on.

OpenStudy (anonymous):

probabaly jus different naming...cos we do have enough tools...so its probably same calc....my calc 2 myt be ur 3,bcos i have calc 1 then calc 2,then adv calc 1 then adv calc 2....wats @Loser66 Major

OpenStudy (anonymous):

Yeah, I know some places even have a calc 4. If I follow the university math path, its calc 1, calc 2, discrete, calc 3, linear algebra, ODE

OpenStudy (anonymous):

Not really, but we had to have calc 2 first. Was a horrible class. So poorly organized, I got a B in it and felt like I learned nothing.

OpenStudy (anonymous):

Well, I was talkign about discrete, lol.

OpenStudy (anonymous):

B in discrete because it just was a stupid class.

OpenStudy (anonymous):

It is, but the class I was in was stupid. Poorly taught course.

OpenStudy (anonymous):

Bingo. She was so out of it, unorganized, sped through lectures, explained things poorly, it was a waste. I just figure itll be better for discrete 2.

OpenStudy (anonymous):

Eh, I just was disappointed with it is all. Now just waiting for the semester to start up again. Oh, well I go back and forth between the community college and the university here. Community college is College of Southern Nevada, uni is UNLV

OpenStudy (anonymous):

But I need to head out, been keeping family waiting, lol. Thanks again and late Merry Christmas.